If m = 2, find the value of:
m – 2 …..................

Grade:12th pass

2 Answers

Harshit Singh
2 years ago
Dear Student

From the question it is given that m = 2
Then, substitute the value of p in the question
= 2 -2
=0

Thanks
Shraddha Singh
2 years ago
Dear Student, 
Given : m= 2
Substituting value of m in equation,
 
m - 2 = 2 - 2 = 0
 
Hope it helps!
Thank you ✌️